Find the critical speed of the ball mill in RPM?


(a) 20.1


(b) 30.7 (c) 49.6 (J) 67.8. (e) None


1
Expert's answer
2021-09-04T01:01:20-0400

Nc = 1/2π[√g/(R-r)]--------(1)

Where Nc is the critical speed of the ball mill in revolution per second

g= acceleration due to gravity= 9.8 m/s2

Consider

R= Radius of ball mill in m =0.9m

r= radius of ball in m=0.03 m

Put all values in above equation

Nc = 1/2π[√g/(R-r)]

Nc=1/2×3.14[√9.81/(0.9-0.03)]

Nc= 1/6.28[√9.81/0.87]

=1/6.28[√11.27]

= 0.534rps

= 0.534×60rpm

Nc=32.04 rpm

Answer is none of the above